Suppose that a pharmacist has a supply of a 30% solution of alcohol and a 70% solution of alcohol. How many quarts of each should be mixed to produce 60 quarts that is 50% alcohol?
thank you

Answer by mananth(16946)

---------percent ---------percent ---------------- quantity
Alcohol I 30 ---------------- x quarts
Alcohol II 70 ---------------- 60 - x quarts
Mixture 50.00% ---------------- 60

30 x + 70 ( 60 - x ) = 50 * 60

30x+4200-70x=3000
30x-70x =3000 -4200
-40 x = -1200
/ -40
x = 30 quarts 30.00% Alcohol I
30 quarts 70.00% Alcohol II